HW09: Problem 6 Previous Problem List Next (4 points) A tire manufacturer produces tires that have a mean life of at least 27500 miles when the production process is working properly. The operations manager stops the production process if there is evidence that the mean tire life is below 27500 miles. The testable hypotheses in this situation are Ho: H 27500 vs HA H27500 1. Identify the consequences of making a Type I error. A. The manager stops production when it is not necessary. B. The manager stops production when it is necessary. C. The manager does not stop production when it is not necessary. D. The manager does not stop production when it is necessary. 2. Identify the consequences of making a Type ll error A. The manager does not stop production when it is necessary. B. The manager stops production when it is necessary. C. The manager stops production when it is not necessary. D. The manager does not stop production when it is not necessary To monitor the production process, the operations manager takes a random sample of 35 tires each week and subjects them to destructive testing. They calculate the mean life of the tires in the sample, and if it is less than 26750, they will stop production and recalibrate the machines. They know based on past experience that the standard deviation of the tire life is 2500 miles. 3. What is the probability that the manager will make a Type I error using this decision rule? Round your answer to four decimal places. 4. Using this decision rule, what is the power of the test if the actual mean life of the tires is 26600 miles? That is, what is the probability they will reject Ho when the actual average life of the tires is 26600 miles? Round your answer to four decimal places.
Answer #1

1) option A is correct

2)option A is correct

3)

P(type i error)=P(Xbar<26750 given mean=27500)=P(Z<(26750-27500)*sqrt(35)/2500)=P(Z<-1.77)=0.0384 ( please try 0.0379 if this comes wrong and revert)

4)

P(power)=P(Xbar<26750 given mean=26600)=P(Z<(26750-26600)*sqrt(35)/2500)=P(Z<0.35)=0.6368 ( please try 0.6387 if this comes wrong and revert)